OpenStudy (anonymous):

A right triangle has a hypotenuse of 12.5 feet and a side of 7.5 feet. What is the measure of the third side?

OpenStudy (anonymous):

pythagorean theorem : a^2 + b^2 = c^2 7.5^2 + b^2 = 12.5^2 56.25 + b^2 = 156.25 b^2 = 156.25 - 56.25 b^2 = 100 b = sq rt 100 b = 10 check.. 7.5^2 + 10^2 = 12.5^2 56.25 + 100 = 156.25 156.25 = 156.25 (correct) the missing side is 10 feet